OpenStudy (anonymous):

solve. check for extraneous solutions. sqrt x+56=x 4 -4 8 -8

OpenStudy (mathstudent55):

\(\sqrt{x} + 56 = x\) Move the 56 to the right side by subtracting it from both sides. Then square both sides to get rid of the square root. What do you get?

OpenStudy (anonymous):

x^2-x-56=0 x^2+(a+b)x+ab=(x+a)(x+b) (x-8)(x+7)=0 x-8=0 x=8 x+7=0 x=-7 x=8,-7

OpenStudy (mathstudent55):

@E.ali You solved your equation correctly, but that equation is not equivalent to the equation in the original post. To find the equivalent equation without a square root, follow my suggestion above. Then make sure you check the solutions in the original equation since squaring both sides of an equation may introduce extraneous solutions.